Ali’s budget line relating good X and good Y has intercept of 50 unit of good X and 20 units of good Y. if the price of good X is 12, what is Ali’s income? What is the price of good Y? What is then slope of budget line?


1
Expert's answer
2021-08-16T08:57:50-0400

If a budget line relating good X and good Y has intercept Qx = 20 unit of good X and Qy = 30 units of good Y and Py = $10, then the Ali’s income is:

I = Px*Qx + Py*Qy

When the budget line has the intercept Qy = 30 and Py = $10, then I = Px*0 + 30*10 = 300, so the Ali's income is I = $300.

The price of good Y can be found from the income equation in the point of interception with the X-axis:

300 = Px*20 + 10*0,

Px = $15.

The slope of budget line is:

300 = 15*Qx + 10*Qy

10Qy = 300 - 15Qx

Qy = -1.5Qx + 30, so the slope of budget line is -1.5.